OpenStudy (anonymous):

Which options could be the next step in simplifying this expression? 5 • (2 + 4) Choose exactly two answers that are correct. A. 5 • 2 + 4 (use the distributive property) B. 5 • 2 + 5 • 4 (use the distributive property) C. 5 • 6 (use order of operations) D. (5 • 2) + 4 (use the associative property)

OpenStudy (anonymous):

@sammixboo @CrashOnce @DeltaSquad @MARC_

sammixboo (sammixboo):

What do you think?

OpenStudy (anonymous):

d b ?

sammixboo (sammixboo):

B is correct because 8(2+4) = 8*6 = 48 8*2 + 8*4 = 16 + 32 = 48 D is not correct though, because 8(2+4) = 8*6 = 48 (8*2) + 4 = 16 + 4 = 20

OpenStudy (anonymous):

so b c

sammixboo (sammixboo):

They don't equal the same, but how do you start solving 5(2+4) ?

sammixboo (sammixboo):

Correct

OpenStudy (anonymous):

Martha biked quite a few days last month. She biked 25 miles the first day and 15 miles each day after that. Let m represent the number of days after the first day. Which expression shows the total distance Martha biked last month? A. (15 + 25)m B. 25m + 15 C. 15m + 25 D. 15 + 25 + m

sammixboo (sammixboo):

What do you think? if she rides on her bike 15 miles every day for the next m days you are multiplying 15m

OpenStudy (anonymous):

c

sammixboo (sammixboo):

And then she rode her bike for 25 miles the first day, so you add 25 and you get 15m + 25, which is C

sammixboo (sammixboo):

So you are correct

OpenStudy (anonymous):

Tomas bought some fruit at the market. He bought twice as many melons as apples. He bought 12 more melons than oranges. He bought 11 apples. How many pieces of fruit did Tomas buy

sammixboo (sammixboo):

Hey can you post your questions on separate threads so we don't spam other user's notifications :P

OpenStudy (anonymous):

ok